# Community detection for Contexual-LSBM: Theoretical limitation on misclassfication ratio and effecient algorithm

### v1

@@ 3.2 Lower bound on number of misclassfication nodes
$D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A}):=\max\left\{\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k),\mathbf{P}(k,k_{1})\right),\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k),\mathbf{P}(k,k_{2})\right)\right\},$
@@
$D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X}):=\max\biggl{\{}\operatorname{KL}_{X}\left(q_{X},p_{X}(k_{1})\right),\operatorname{KL}_{X}\left(q_{X},p_{X}(k_{2})\right)\biggr{\}}.$ (5)
[…]
Theorem 1.
Denote $\bar{p}:=\max_{i,j,l\geq 1}\mathbf{P}(i,j,l)$, grant $\bar{p}=\omega(1/n)$, $\bar{p}=o(1)$ and 1. One has
lim $\liminf\frac{nD(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})}{\log(n/s)}\geq 1$ (6)
if $\eta_{2}=o(n)$. Here we assume $\limsup_{n\rightarrow\infty}\frac{\bar{s}}{s}\leq 1$. Here $\bar{s}$ is defined in (2).
@@
See Appendix A. ∎
Theorem 1 indeed implies that the number of misclassfied nodes for the proposed CLSBM model is above the order $n\exp(-nD(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu}))$. The following remark indeed builds the connections between the Chernoff information $I$ proposed in (Dreveton et al., 2024, equation $3.4$) and $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ in (4)
@@
$=\min_{k_{1}\neq k_{2}}\max_{t\in[0,1]}\sum_{k}^{K}\alpha_{k}\biggl{[}\sum_{l=1}^{L}\left((1-t)\mathbf{P}(k,k_{1},l)+t\mathbf{P}(k,k_{2},l)-\mathbf{P}^{1-t}(k,k_{1},l)\mathbf{P}^{t}(k,k_{2},l)\right)+\frac{t(1-t)}{2n}\left\|\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\right\|_{2}^{2}\biggr{]}.$ (7)
Proof.
See Appendix B. ∎
From Remark 1, if the contextual attribute is missing such that $\boldsymbol{\mu}=0$, the divergence $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ reduces to the CH-divergence, as stated in Yun and Proutiere (2016). On the other hand, when the topological information is missing (e.g., $\mathbf{P}(k,k_{1},l)=\mathbf{P}(k,k_{2},l)$ for all $k,k_{1},k_{2},l$), we have $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})=\frac{1}{8n}\|\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\|_{2}^{2}$, which corresponds to the Gaussian mixture model, as shown in (Lu and Zhou, 2016, Theorem 3.3).
3.3 Spectral clustering for latent factor model
In this section, we extend the proposed CLSBM to a semiparametric model and demonstrate that the spectral method can consistently estimate the proposed estimator $s(\hat{\sigma})$ defined in (10), given the adjacency matrices $\{\mathbf{A}_{l}\}_{l\in\{0,1,2,\dots,L\}}$ and node attribute vectors $\{\mathbf{x}_{i}\}_{i\in[n]}$. Specifically, we first outline the construction of the estimator by combining the topological information $\{\mathbf{A}_{l}\}_{l\in\{0,1,2,\dots,L\}}$ with a Gaussian mixture model and provide an intuitive explanation of the spectral method at the population level in Section 3.3.1. Subsequently, in the following section, we formalize the spectral clustering method in Section 3.3.2 and establish the consistency of the proposed method, as proven in Lemma 1, to demonstrate its effectiveness.
@@ 3.3.1 Aggregated latent factor model
The latent factor model demonstrates its capability in revealing hidden low-dimensional structures within observed high-dimensional data and has been widely applied in various domains, including PCA, Gaussian mixtures, Nonnegative Matrix Factorization, and Sparse Factor Analysis. This provides an intuitive basis for combining both the topological model and Gaussian mixture into a shared latent space, as both the adjacency matrix and the observed Gaussian mixture share the same column space spanned by the ground truth assignment matrix $\mathbf{Z}(\sigma)\in\mathbb{R}^{n\times k}$, where $\mathbf{z}_{i\cdot}:=\mathbf{e}_{\sigma(i)}^{T}$. To elaborate, we begin by defining our aggregator factor matrix $\mathbf{S}$: given weights $\{w_{l}\}_{l\in[L]}$ independently and identically sampled from a uniform distribution over $[0,1]$, we define
$\mathbf{S}:=\sum_{l=1}^{L}w_{l}\mathbf{A}^{(l)}+\frac{1}{n}\mathbf{X}\mathbf{X}^{T}.$ (8)
where $\mathbf{X}\in\mathbb{R}^{n\times d}$ represents the collection of $n$ observed Gaussian mixture data. The weights $\{w_{l}\}_{l\in[L]}$ are used to identify signals for different labels $l\in[L]$, and the factor $1/n$ balances the signal-to-noise ratio (SNR) between the graph signal and the Gaussian signal, as motivated by Remark 1. Let $\mathbf{P}_{s}:=\frac{1}{L}\sum_{l=1}^{L}\mathbf{P}_{\cdot,\cdot,l}\in\mathbb{R}^{K\times K}$. At the population level, we have:
@@
$=\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T}-\text{diag}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T})+\frac{1}{n}\mathbf{Z}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\mathbf{Z}^{T}$
$=\underbrace{\mathbf{Z}\left(\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\right)\mathbf{Z}^{T}}_{\mathbf{M}}-\text{diag}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T}).$ (9)
Here, $\mathcal{H}(\cdot)$ denotes the hollowing operation, which removes the diagonal elements of the matrix, and $\mathbf{Z}$ is shorthand for $\mathbf{Z}(\sigma)$. It is evident that the first term $\mathbf{M}$ in $\mathbb{E}\left[\mathbf{S}\right]$ is a symmetric low-rank factor model, which has been extensively studied in areas such as low-rank matrix decomposition, Factor Analysis, signal processing, and spectral clustering. By the definition of $\mathbf{M}$, it is clear that the optimal solution of the $k$-means algorithm applied to the columns/rows of $\mathbf{M}$ should recover the ground truth assignment $\mathbf{Z}$. In practice, however, instead of directly applying $k$-means to the columns of $\mathbf{M}$ (which has dimension $n$), spectral methods are often employed due to their computational advantages. Specifically, let $\mathbf{U}_{K}$ denote the top $K$ eigenvectors of $\mathbf{M}$. We apply $k$-means to the columns of $\mathbf{U}_{K}^{T}\mathbf{M}$. This is equivalent to applying $k$-means to the columns of $\mathbf{M}$ or $\mathbf{U}_{K}^{T}\mathbf{M}$ in terms of clustering outcomes, as $\mathbf{U}_{K}$ is an orthonormal matrix that preserves distances during projection. However, since neither $\mathbf{M}$ nor $\mathbf{U}_{K}$ is observed, even at the population level, we substitute $\mathbf{S}$ for $\mathbf{M}$ and use its top eigenvectors, denoted as $\bar{\mathbf{U}_{K}}$, in place of $\mathbf{U}_{K}$. Given that the second term in (3.3.1) is small in operator norm, i.e., $\|\text{diag}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T})\|_{\text{op}}=o(1)$, the misclassification rate of the $k$-means algorithm based on $\mathbf{S}$ depends on $\|\mathbf{S}-\mathbb{E}(\mathbf{S})\|_{F}$, which can be bounded using concentration inequalities. We summarize the spectral clustering procedure in Algorithm 1.
1:Node attributes $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$, adjacent matrix $\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}}$, and number of clusters $K$
2:Cluster assignments for each data point
@@
$\bar{\sigma}(\cdot),\left\{\bar{\boldsymbol{\theta}}\right\}_{i\in[K]}=\operatorname*{argmin}_{\begin{subarray}{c}\sigma\in\mathcal{F}_{a},\\ \boldsymbol{\theta}_{i}\in\mathbb{R}^{k}\end{subarray}}\sum_{i\in[n]}\left\|\mathbf{Q}_{i}-\boldsymbol{\theta}_{\sigma(i)}\right\|_{2}^{2}.$
7:Return Cluster assignments function $\bar{\sigma}(\cdot)$.
Algorithm 1 Spectral Clustering Algorithm
@@ 3.3.2 Theoretical guarantee
Denote $s(\bar{\sigma})$ as the number of misclassified nodes for label function $\bar{\sigma}$ recovered from Algorithm 1 such that
$s(\bar{\sigma}):=\inf_{\pi\in S_{n}}~{}\mathbb{E}\left[\sum_{i\in n}\mathbf{1}_{\{(\bar{\sigma}(\pi(i)))\neq\sigma(i))\}}\right].$ (10)
@@
Grant the assumptions in theorem 1 and further assume $\bar{p}=w(\log(n)/n)$, let $\bar{\sigma}$ be the solution of the spectral algorithm, we have
$s(\bar{\sigma})\lesssim\frac{K}{n\textit{SNR}}$ (11)
with high probability. Here $\textit{SNR}:=(1+o(1))\min_{k_{1}\neq k_{2}}\left\|\left(\mathbf{P}_{s}\right)_{k_{1}}-\left(\mathbf{P}_{s}\right)_{k_{2}}+\frac{1}{n}\left(\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\right)^{T}\boldsymbol{\mu}\right\|_{2}^{2}$. Furthermore if $n\textit{SNR}=w(1)$ we obtain $s(\bar{\sigma})=o(1)$.
@@
See Appendix C
Lemma 10 demonstrates that the proposed indicator $s(\bar{\sigma})$ is upper bounded by the inverse of the defined (SNR), which corresponds to the minimum distance between any two columns in the matrix $\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}$. Rates of this nature have been extensively studied in the literature Lei and Rinaldo (2015); Löffler et al. (2021); Zhang and Zhou (2024) for spectral methods under SBM or Gaussian mixtures, respectively. However, our work is the first to provide theoretical guarantees for spectral methods under our proposed CLSBM. Despite achieving consistency, our analysis does not attain the optimal rate presented in Theorem 1, which is exponentially small. However, we beileive that this estimate could indeed play the role of a initializers for the potential refinement as seen in Yun and Proutiere (2016); Gao et al. (2018).
@@ Appendix A Proof of Theorem 1
Mimic the change of measure method in Yun and Proutiere (2016) we adopt it to the CLSBM model and its pertubted version in this section. Denote $\Phi$ and $\Psi$ as the parameter for the true parameter model(CLSBM) and one potential perturbed parameter model respectively. To be more specific, the true ground truth model $\Phi$ first generate the cluster for each node based on $\boldsymbol{\alpha}$ such that $\mathbb{P}(\sigma(i)=k)=\alpha_{k}$ for the $i$-th node indepedently. For the topological imformation, model $\Phi$ generate the observed label $l\in[L]$ for two different node $i$, $j$ based on the $\mathbf{P}(\sigma(i),\sigma(j),l)$. For the contexual information, $\mathbf{x}_{i}$ is sampled from $\mathcal{N}(\boldsymbol{\mu}_{\sigma(i)},\mathbf{I}_{d})$. We state the definition of $\Psi$ as follows:
@@
The edge label generation mechanism for the topological information($\mathbf{A}$) in $\Psi$ is the same as in Yun and Proutiere (2016) by replace $\mathbf{q}$ with $\mathbf{q}_{\mathbf{A}}^{\star}$, here we state the generation of label for side information $\{\mathbf{x}_{i}\}_{i\in[n]}$ as follows: denote $i^{\star}=\operatorname*{argmin}_{\sigma(i)\in\{k_{1}^{\star},k_{2}^{\star}\}}i$.
  • When $i\neq i^{\star}$: $\mathbf{x}_{i}$ is sampled from $\mathcal{N}(\boldsymbol{\mu}_{\sigma(i)},\mathbf{I}_{d})$ which is same as in $\Phi$.
@@
We denote the $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ as the pseudolikelihood ratio of the observed labels and contexual variables such that
@@
In the following two sections we thus provide the detailed proof of theorem 1, which follows the proof in Yun and Proutiere (2016). The proof consist of two parts, the first part constructs a stochastic model $\Psi$ and lower bound mis-classification node $\mathbb{E}_{\Phi}(\epsilon^{\pi}(n))$ by the expectation of the proposed log-likelihood ratio $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ defined in (13); In the second part we further estimate the ratio upper bound of $\mathbb{E}_{\Psi}(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}})$ and $\mathbb{E}_{\Psi}\left[\left(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}-\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]\right)^{2}\right]$ respectively to finish the proof.

### v3

@@ 3.2 Lower bound on the number of misclassification nodes
$D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A}):=\max\biggl{\{}\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k)~{}\|~{}\mathbf{P}(k,k_{1})\right),\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k)~{}\|~{}\mathbf{P}(k,k_{2})\right)\biggr{\}},$
$D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X}):=\max\biggl{\{}\operatorname{KL}_{X}\left(q_{X}~{}\|~{}p_{X}(k_{1})\right),\operatorname{KL}_{X}\left(q_{X}~{}\|~{}p_{X}(k_{2})\right)\biggr{\}}.$ (5)
@@
Theorem 1 (Lower bound).
Denote $\bar{p}:=\max_{i,j,l\geq 1}\mathbf{P}(i,j,l)$, grant 1 and assume $\bar{p}=\omega(1/n)$, $\bar{p}=o(1)$, and $\eta_{2}=o(n)$. Let $s=o(n)$. If there exists an algorithm that asymptotically has fewer misclassified nodes than $s$ in expectation, i.e., $\limsup_{n\rightarrow\infty}\frac{\bar{s}}{s}\leq 1$, then we have
$\liminf_{n\rightarrow\infty}\frac{nD(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})}{\log(n/s)}\geq 1,$ (6)
where $\bar{s}$ is defined in (2).
@@
We defer the proof to Appendix A. ∎
Theorem 1 implies that the expected number of misclassified nodes for the proposed CLSBM is at least of the order $n\exp(-nD(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu}))$. The following remark indicates that the proposed measure $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ generalizes both the CH-divergence for the LSBM (Yun and Proutiere, 2016, Claim 4) and the minimal distance between the centers of Gaussian mixtures.
@@
$\!\!=\min_{k_{1}\neq k_{2}\in[K]}\max_{t\in[0,1]}\sum_{k=1}^{K}\alpha_{k}\biggl{[}\sum_{l=1}^{L}\biggl{(}(1-t)\mathbf{P}(k,k_{1},l)+t\mathbf{P}(k,k_{2},l)-\mathbf{P}^{1-t}(k,k_{1},l)\mathbf{P}^{t}(k,k_{2},l)\biggr{)}\!\!+\frac{t(1-t)}{2n}\left\|\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\right\|_{2}^{2}\biggr{]}.$ (7)
The proof is deferred Appendix B. When the contextual node attributes are absent (i.e., $\boldsymbol{\mu}=0$), the divergence $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ simplifies to the CH-divergence introduced in Yun and Proutiere (2016). On the other hand, in cases where topological information is missing (e.g., $\mathbf{P}(k,k_{1},l)=\mathbf{P}(k,k_{2},l)$ for all $k,k_{1},k_{2},l$), we have $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})=\frac{1}{8n}\|\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\|_{2}^{2}$, which corresponds to threshold for the Gaussian mixture model, as shown in (Lu and Zhou, 2016, Theorem 3.3).
3.3 Spectral community detection for latent factor model
In this section, we extend the proposed CLSBM to a latent factor model and demonstrate that the spectral-based method can consistently estimate the proposed estimator $s(\hat{\sigma})$ defined in (10), based on the adjacency matrices $\{\mathbf{A}_{l}\}_{l\in\{0,1,2,\dots,L\}}$ and contextual node attribute vectors $\{\mathbf{x}_{i}\}_{i\in[n]}$. Specifically, we first outline the construction of the estimator by combining the topological information with the contextual node attributes and provide an intuitive explanation of the spectral-based method in Section 3.3.1. Subsequently, in the following section, we formalize the spectral-based method in Section 3.3.2 and establish its consistency, as proven in Lemma 1, thereby demonstrating its effectiveness.
@@ 3.3.1 Aggregated latent factor model
The latent factor model demonstrates its capability in revealing hidden low-dimensional structures within observed high-dimensional data and has been widely applied in various domains, including PCA, Gaussian mixtures, nonnegative matrix factorization, sparse factor analysis and so on. This provides an intuitive basis for combining both the topological model and contextual node attributes into a shared latent space, as both the adjacency matrix and the node attributes share the same eigenspace spanned by the ground truth assignment matrix $\mathbf{Z}(\sigma)\in\mathbb{R}^{n\times k}$, where $\mathbf{Z}_{i\cdot}:=\mathbf{e}_{\sigma(i)}^{T}$. To elaborate, we begin by defining our aggregator factor matrix $\mathbf{S}$: Given weights $\{w_{l}\}_{l\in[L]}$ i.i.d sampled from a uniform distribution over $[0,1]$, we define:
$\mathbf{S}:=\sum_{l=1}^{L}w_{l}\mathbf{A}_{l}+\frac{1}{n}\mathbf{X}^{T}\mathbf{X},$ (8)
where $\mathbf{X}\in\mathbb{R}^{d\times n}$ is a matrix whose columns represent the collection of $n$ observed node attributes. The weights $\{w_{l}\}_{l\in[L]}$ are employed to identify signals for different labels $l\in[L]$, and the factor $1/n$ balances the signal-to-noise ratio (SNR) between the graph signal and the Gaussian signal, as motivated by Remark 1. Let $\mathbf{P}_{s}:=1/(2L)\sum_{l=1}^{L}\mathbf{P}(\cdot,\cdot,l)\in\mathbb{R}^{K\times K}$. At the population level, the following holds:
@@
$=\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T}-\operatorname{diag}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T})+\frac{1}{n}\mathbf{Z}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\mathbf{Z}^{T}$
$=\underbrace{\mathbf{Z}\left(\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\right)\mathbf{Z}^{T}}_{\mathbf{M}}-\operatorname{diag}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T}).$ (9)
Here, $\mathcal{H}(\cdot)$ denotes the hollowing operation, which sets the diagonal elements of the matrix to zero, and $\mathbf{Z}$ is shorthand for $\mathbf{Z}(\sigma)$. It is evident that the first term $\mathbf{M}$ in $\mathbb{E}\left[\mathbf{S}\right]$ is a symmetric low-rank factor model when $K<n$, which has been extensively studied in areas such as low-rank matrix decomposition, factor analysis, signal processing, and spectral clustering. By definition, the optimal solution of the $k$-means algorithm applied to the columns/rows of $\mathbf{M}$ should yeild the ground truth assignment $\mathbf{Z}$. In practice, however, instead of directly applying $k$-means to the columns of $\mathbf{M}$ (which has dimension $n$), spectral methods Newman (2006) are often pre-employed due to their computational effficiency. Specifically, let $\mathbf{U}_{K}\in\mathbb{R}^{n\times K}$ denote the top-$K$ eigenvectors of $\mathbf{M}$. We apply $k$-means to the columns of $\mathbf{U}_{K}^{T}\mathbf{M}$. In fact applying $k$-means to the columns of $\mathbf{M}$ or $\mathbf{U}_{K}^{T}\mathbf{M}$ yeilds equivalent clustering outcomes, as $\mathbf{U}_{K}$ is an orthonormal matrix that preserves distances during projection. Chooseing $\mathbf{U}_{K}^{T}\mathbf{M}$ is motivated by its computational efficiency for clustering, particularly because $K$ is typically much smaller than $n$. However, since neither $\mathbf{M}$ nor $\mathbf{U}_{K}$ is observed even at the population level, we substitute $\mathbf{S}$ for $\mathbf{M}$ and employ its top-$K$ eigenvectors, denoted as $\bar{\mathbf{U}}_{K}$, in place of $\mathbf{U}_{K}$. Given that the second term in (3.3.1) has a small operator norm, i.e., $\|\text{diag}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T})\|_{\text{op}}=o(1)$, the misclassification rate of the $k$-means algorithm based on $\mathbf{S}$ depends on $\|\mathbf{S}-\mathbb{E}(\mathbf{S})\|_{F}$, a quantity can be further bounded via some concentration inequalities. We summarize the aforementioned procedures in Algorithm 1, refer to as Spectral Community Detection.
1:Node attributes $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$, adjacency matrices $\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}}$, and number of communities $K$.
2:Assignment for each node
@@
$\bar{\sigma}(\cdot),\left\{\bar{\boldsymbol{\theta}}\right\}_{i\in[K]}=\operatorname*{argmin}_{\begin{subarray}{c}\sigma\in\mathcal{F}_{a},\\ \boldsymbol{\theta}_{i}\in\mathbb{R}^{k}\end{subarray}}\sum_{i\in[n]}\left\|\mathbf{Q}_{\cdot i}-\boldsymbol{\theta}_{\sigma(i)}\right\|_{2}^{2}.$
7:Return: Assignment indicator function $\bar{\sigma}(\cdot)$.
Algorithm 1 Spectral Community Detection
@@ 3.3.2 Theoretical guarantee
Denote $s(\bar{\sigma})$ as the number of misclassified nodes for indicator function $\bar{\sigma}$ recovered from Algorithm 1 such that
$s(\bar{\sigma}):=\inf_{\pi\in S_{K}}~{}\mathbb{E}\left[\sum_{i\in n}\mathbf{1}_{\{\pi(\bar{\sigma}(i))\neq\sigma(i))\}}\right].$ (10)
@@
Grant the assumptions in Theorem 1 and further assume $\bar{p}=\omega(\log(n)/n)$, let $\bar{\sigma}$ be the solution of the Algorithm 1, we have
$s(\bar{\sigma})\lesssim\frac{K}{n\cdot\mathrm{SNR}}$ (11)
with high probability, where $\mathrm{SNR}:=\min_{k_{1}\neq k_{2}}\left\|\left(\mathbf{P}_{s}\right)_{k_{1}}-\left(\mathbf{P}_{s}\right)_{k_{2}}+\frac{1}{n}\left(\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\right)^{T}\boldsymbol{\mu}\right\|_{2}^{2}$. Furthermore if $n\cdot\mathrm{SNR}=w(1)$ we obtain $s(\bar{\sigma})=o(1)$.
@@
The proof is deferred to Appendix C. ∎
Lemma 1 demonstrates that the proposed indicator $s(\bar{\sigma})$ is upper bounded by the inverse of the defined (SNR) in a polynomial manner, which corresponds to the minimum distance between any two columns in the matrix $\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}$. Rates of this nature have been extensively studied in the literature Lei and Rinaldo (2015); Löffler et al. (2021); Zhang and Zhou (2024) for spectral-based methods under SBM or Gaussian mixtures, respectively. However, our work is the first to provide theoretical guarantees for spectral-based methods under the CLSBM. Despite achieving consistency, our analysis does not lead to the optimal rate presented in Theorem 1, which is exponentially small. However, we believe that this estimate can serve as an effective initializer for potential subsequent refinement steps, as demonstrated in Yun and Proutiere (2016); Gao et al. (2018). By leveraging the spectral method’s output as a starting point, further refinement techniques such as MLE can be applied to improve the accuracy of community detection, potentially bridging the gap between polynomial and exponential misclassification rates.
@@ Appendix A Proof of Theorem 1
Mimicking the change of measure method in Yun and Proutiere (2016) we adapt it to the CLSBM model and its perturbed version in this section. Denote $\Phi$ and $\Psi$ as the parameter for the true parameter model (CLSBM) and a potential perturbed parameter model respectively. To be more specific, the true ground truth model $\Phi$ first generates the cluster for each node based on $\boldsymbol{\alpha}$ such that $\mathbb{P}(\sigma(i)=k)=\alpha_{k}$ for the $i$-th node indepedently. For the topological information, model $\Phi$ generates the observed labels $l\in[L]$ for two different nodes $i$, $j$ based on the $\mathbf{P}(\sigma(i),\sigma(j),l)$. For the contextual information, $\mathbf{x}_{i}$ is sampled from $\mathcal{N}(\boldsymbol{\mu}_{\sigma(i)},\mathbf{I}_{d})$. We state the definition of $\Psi$ as follows:
@@
The edge label generation mechanism for the topological information($\mathbf{A}$) in $\Psi$ is the same as in Yun and Proutiere (2016) with replace $\mathbf{q}$ with $\mathbf{q}_{\mathbf{A}}^{\star}$, Here, we describe the generation of labels for the side information $\{\mathbf{x}_{i}\}_{i\in[n]}$ as follows: let $i^{\star}=\operatorname*{argmin}_{\sigma(i)\in\{k_{1}^{\star},k_{2}^{\star}\}}i$.
  • When $i\neq i^{\star}$: $\mathbf{x}_{i}$ is sampled from $\mathcal{N}(\boldsymbol{\mu}_{\sigma(i)},\mathbf{I}_{d})$ which is the same as in $\Phi$.
@@
We denote $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ as the pseudolikelihood ratio of the observed labels and contextual variables such that:
@@
In the following two sections, we provide the detailed proof of theorem 1, which follows the proof in Yun and Proutiere (2016). The proof consists of two parts, the first part constructs a stochastic model $\Psi$ and lower bounds the misclassification error $\mathbb{E}_{\Phi}(\epsilon^{\pi}(n))$ using the expectation of the proposed log-likelihood ratio $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ defined in (13); In the second part, we further estimate upper bounds for $\mathbb{E}_{\Psi}(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}})$ and $\mathbb{E}_{\Psi}\left[\left(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}-\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]\right)^{2}\right]$ respectively, to complete the proof.